MODS1 Commissioning Run 7: UTC 2011 April 14

Observer(s): R. Pogge, R. Stoll (OSU), E. Skillman (UMinn), M. Fumana (INAF), M. Wagner (OSU/LBTO)
Telescope Operator: G. Bechetti (LBTO)
AO Support: Remote from Tucson as needed
Telescope Support: J. Little (LBTO)
Instrument Support: D. Officer (LBTO)

Summary

Clear and cold, but no observing because of high winds all night long

Preparations

Andrew reports that our mirror limits problem was caused by a typographical error that got into the revised collimation model coefficients. We should not have problems tonight.

No special preparation today except routine filling of the dewars.

All 6 mounting bolts checked in the afternoon and all were tight.

Details

All times are UTC unless otherwise noted.

0500 - High winds at sunset, we are waiting them out. As of 0500 we are seeing gusts up to 25m/s.

1150 - High winds continued all night. It was clear, and beautifully dark (if cold, -5°C, and windy) after moonset.

1200 - Called the night officially.

-- RichardPogge - 13 Apr 2011
Topic revision: r3 - 14 Apr 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback